# Just-in-time and distributed task representations in language models

### v2

@@ 3.3 Task representations exhibit scope locality
We have seen evidence that transferrable task representations tend to be temporally local. That leads to the question of whether they have a lasting effect over generation. That is, are the restored task contexts in the zero-shot forward pass also fleeting in nature? Prior work has mostly focused on simple, single-token output tasks. To study the scope of task representations, we tested to what extent restored task contexts can support longer generation beyond the first token. Building on the simple tasks from prior work, we evaluated models on a set of longer-generation and mixed-generation tasks, including repeating a simple task multiple times on different input words, list-level tasks that operate over multiple words, and inferring/performing different tasks on different words (see Methods and Appendix A).
@@
Figure 5: Reinstantiated task contexts in longer- and mixed-generation tasks often decay over generation, especially for tasks that can be decomposed into semantically-independent subtasks. This suggests a tendency for models to only activate transferrable representations for small task scopes. A. Bar plot: recontextualized zero-shot accuracy compared to zero-shot and 8-shot accuracy on longer-generation tasks. Line plots: recontextualization accuracy for each output unit across models, conditioned on sequences where models generated full correct responses with eight examples in-context. An output unit usually corresponds to a single word and is occasionally a short phrase (e.g. the capital of a country). B. Visualization as in A, but for mixed-generation tasks.
Transferrable task representations tend to support a semantically-independent task scope.
@@
In these experiments, we find further evidence of the locality of transferrable task representations. Overall, the recontextualized zero-shot accuracy of tasks that require longer and mixed answers is substantially lower than that in tasks that require shorter answers (Figure 5, bar plots; also see Figure S1). Across a range of tasks, we find that the recontextualized zero-shot accuracy decreases for each output word (Figure 5, line plots; see similar results on Qwen3 models in Figure S13), suggesting that the restored task contexts “fade” over longer generation.
@@
Models seem to be consistently decomposing tasks and forming local task representations that capture a minimal “task scope”. For example, in choose_first_middle_last_of_5, the reinstantiated task context only supported generating the first word, suggesting that further token generations rely on additional representations instantiated by a separate mechanism while processing the previously-generated tokens. This effect is very pronounced in the mixed-generation tasks. In these cases, all models form strong local task contexts that only encapsulate the first subtask in a multi-task chain, and defer representing later subtasks. Interestingly, the extracted task representations for the same simple task when it appears independently or as a first task in a repeated or mixed-task context can be distinct (Figure 4B), even though they support generating the same responses. This can potentially reflect an inert sensitivity to the broader mixed-task context based on vocabulary differences. Additionally, these differences may also convey information relevant for later subtasks to be activated, even though they are not sufficient to directly restore execution of later subtasks.
The restored task contexts do support longer generation to some extent in many tasks, including repeating a simple task multiple times and some list-level tasks (i.e., reversing or shifting a word list). Overall, these results suggest that language models may be automatically segmenting semantically-independent task scopes when possible, such that task representations for longer or composite tasks are offloaded onto multiple tokens.
4 Related work
Since the discovery that large language models exhibit emergent in-context learning (Brown et al.,, 2020), there has been substantial interest in investigating this capability and its mechanistic basis.
From a behavioral perspective, many subsequent works have explored how ICL could develop from implicit meta-learning of data properties (Xie et al.,, 2022; Chan et al.,, 2022), and how this may relate to the broader set of language model capabilities (Chen et al.,, 2024; Lampinen et al.,, 2024). Some of this work has focused on the surprising fragility of ICL to subtle prompt changes (e.g. Sclar et al.,, 2024); conversely, others have highlighted how ICL may be overly robust, allowing “learning” common tasks even if the labels are randomized (Min et al.,, 2022). One particularly relevant focus of behavioral work on ICL has been on the dynamics of in-context learning; for example, how adding many example shots can improve performance on difficult tasks, or even those discouraged in post-training (Agarwal et al.,, 2024; Anil et al.,, 2024).
From a mechanistic perspective, Olsson et al., (2022) showed how induction heads could support in-context learning, and other work has studied how they might develop over training (Edelman et al.,, 2024; Singh et al.,, 2025). More recently, attention has turned to cases where models may create internal task representations. These are representations that can be extracted from few-shot prompts and then injected (without the few-shot examples in context) to induce task performance. Hendel et al., (2023) demonstrated an instance of such task representations: representations at intermediate layers of the model at a key token that can be injected to mimic the effect of a few-shot prompt. Concurrent work from Todd et al., (2024) identified “function vectors,” which aggregate the effects of multiple attention heads that convey task information. Subsequent work has generalized and extended these findings, for example exploring how function vectors can emerge from instructions as well as examples (Davidson et al.,, 2025) and how task vectors capture task representations in multimodal models (Huang et al.,, 2024; Luo et al.,, 2024). Other works have explored how these representations and structures emerge over training (Yang et al., 2025b, ; Yin and Steinhardt,, 2025), and extended existing methods to more robustly restore task contexts in zero-shot settings (Li et al.,, 2024). Our work builds on these findings and use transferrable task contexts as a window to study the dynamics of in-context task representations in language models.
@@ 5 Discussion
We sought to understand the dynamics of in-context task representations that support language models’ successful learning of new tasks. Building on prior methods, we evaluated when in the context we can extract transferrable task representations from few-shot settings that restore task contexts in zero-shot settings. Our results show that in many tasks, models refine task representations over more evidence such that the representations more successfully restore task contexts. However, these transferrable task representations only sporadically activate, and seem to best support a minimal task scope (e.g., a first irreducible subtask). The dynamics of these effective, transferrable task representations strongly contrast with a general sensitivity to high-level task differences that persists throughout the context. In addition, we find many cases where models do not appear to condense global task representations into a local, token-level representation, such as tasks involving more state tracking and tasks chaining different types of subtasks together.
In general, our results complicate the intuitive picture that the computational process underlying language models’ in-context learning and task inference is smooth and gradual. Models do not appear to refine task representations at a per-token basis, not even in a step-wise manner—even on a per-example basis, the task state is not sustained, but often fades and reactivates across different tokens. That language models elect to condense task evidence onto a single token at intermediate layers rather than relying on repeated cross-token attention to aggregate information may suggest a general inductive bias to compress knowledge to local representations when possible. This may also relate to the success many works have observed on extracting transferrable task representations in broad settings, such as following instructions or images (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024), or capturing information for multiple possible task outcomes (Xiong et al.,, 2024).
@@
However, one trend that arose from these investigations is that language models do not seem able to condense task information into local representations in all cases. Rather, they exhibit a tendency to form sharp local contexts for small task units, and offload broader task contexts such as mixed or multiple tasks across time. As discussed earlier, even for simple tasks where some intermediate-state tracking may be required, successful inference may need to rely on cross-token and cross-layer computation (e.g., as shown in Ameisen et al.,, 2025). In these cases, the effective restoration of the computation process may also require intervening multiple components during the forward pass. It is also possible that by overriding token activations at intermediate layers with task vectors, the models have lost any task state information formed in earlier layers. Other methods that restore task states through additive injection rather than activation patching may therefore be more successful at restoring model task states in these tasks (Todd et al.,, 2024; Li et al.,, 2024). We also note that different models may differ in the extent to which they form local or distributed effective representations across the context or layers.
@@
An intriguing direction for future work is to study the mechanistic bases for the strong temporal and scope locality we observed in models’ in-context task representations. One possibility may be that the residual stream is more stable and easier to learn from during training. This may encourage the model to rely more on the residual stream to condense contextual task representations rather than relying on the more expensive attention operations. These learning dynamics may drive models to conform with an implicit normative consideration to not instantiate task contexts until needed and instantiate just the right scope to avoid capacity waste. Some of these features may even be exclusive to models pre-trained on natural languages (e.g. Yang et al., 2025b, ).
@@
We note a few important limitations of our work. First, we primarily rely on task vectors as the method to extract transferrable task representations. This means that our conclusions and speculations are bounded by the effectiveness of this method. As we discussed earlier, representations for some task contexts may be more distributed, either across tokens and/or across model layers. It would be important to confirm if similar dynamics are observed in less-constrained methods such as a multi-layer recontextualization (Li et al.,, 2024). Second, we mostly explored relatively simple tasks, including when we investigated longer-generation tasks. It’s possible that many of the dynamics we observe here would not generalize to settings with naturalistic languages, especially when the tasks are not so cleanly decomposable and a single, semantically-independent task unit is hard to define.
@@ Conclusion
We investigated how the dynamics of in-context learning are reflected in the development of language models’ internal task representations. Our results suggest that language models do not smoothly refine a global task state in-context. While general task sensitivity persists throughout context, models appear to construct effective task representations in a “just-in-time” fashion to solve a simple task scope immediately ahead. The fleeting, minimally-scoped nature of these in-context task representations provides new insight into the models’ state of inferring and performing tasks based on new evidence.
@@
We thank Yasaman Bahri, Katherine Hermann, Mike Mozer, Sjoerd van Steenkiste, and Murray Shanahan for helpful comments and support on this work.
@@ Appendix A Tasks
Table 1: Simple/shorter-answer tasks. See Todd et al., (2024) for more details for the first nine tasks.

### v3

@@ 4.3 Transferrable task representations capture variant scope locality
We have seen evidence that transferrable task representations tend to be temporally local. That leads to the question of whether they have a lasting effect over generation. That is, are the restored task contexts in the zero-shot forward pass also fleeting in nature? To study the task scope supported by these representations, we tested to what extent restored contexts can support longer generation beyond the first token. Building on the simple tasks from prior work, we evaluated models on a set of longer-generation and mixed-generation tasks, including repeating a simple task multiple times on different input words, list-level tasks that operate over multiple words, and inferring/performing different tasks on different words (see Methods and Appendix A).
@@
Figure 5: Reinstantiated task contexts in longer- and mixed-generation tasks often decay over generation, especially for tasks that can be decomposed into semantically-independent subtasks. This suggests a tendency for models to only activate transferrable representations for small task scopes. A. Bar plot: recontextualized zero-shot accuracy compared to zero-shot and 8-shot accuracy on longer-generation tasks; accuracies within each task are averaged across output units. Line plots: recontextualization accuracy for each output unit, conditioned on sequences where models generated full correct responses with eight examples in-context. An output unit usually corresponds to a single word and is occasionally a short phrase (e.g. the capital of a country). B. Visualization as in A, but for mixed-generation tasks.
Transferrable task representations sometimes only support limited task scopes.
@@
In these experiments, we find further evidence on the semantic locality of transferrable task representations. Overall, the recontextualized zero-shot accuracy of tasks that require longer and mixed answers is substantially lower than that in tasks that require shorter answers (Figure 5, bar plots; also see Figure S1). Across a range of tasks, we find that the recontextualized zero-shot accuracy decreases over longer generation (Figure 5, line plots; see similar results on Qwen3 models in Figure S13), suggesting that the restored task contexts sometimes “fade” over longer generation.
We notice a strong contrast between different tasks. In some longer-generation tasks that repeat the same task multiple times or require list-level operations, the reinstantiated task contexts (from task vectors extracted from and patched onto the key colon token) successfully supported continued generation. In other tasks, however, this intervention only supported generating the first output. This effect is very pronounced in the mixed-generation tasks that combine multiple distinct subtasks. In these cases, all models form strong local task representations that only encapsulate the first subtask. This suggests that transferrable task representations at the key colon token can capture variant “task scopes”, which tend to be limited in cases where the subtasks have distinct semantic boundaries.
@@
We confirm that models defer representing later subtasks using additional interventions (see Appendix F). Indeed, we find that transferrable subtask representations in mixed-generation tasks form at the comma tokens prior to subtask answer generation. However, intervening comma tokens alone does not seem to recover the full subtask inference ability. We posit that the full subtask representation is distributed across colon and comma tokens. Interestingly, the extracted task representations from the colon tokens in the same simple task can be distinct when it appears independently or as a first task in a multi-task context (Figure 4B), even though they support generating the same responses. This may in-part reflect vocabulary differences in later subtasks, but potentially also information relevant for later subtasks to be activated, even though these representations alone are not sufficient to restore execution of later subtasks.
@@ 5 Discussion
We sought to understand the dynamics of in-context task representations that support language models’ successful learning of new tasks. We evaluated when in the context we can extract transferrable task representations that restore task contexts in zero-shot settings, and when we can extract robust representations of task identity. Our results show that, transferrable task representations only sporadically activate, even though they accrue evidence from multiple examples. However, models maintain a strong general sensitivity to high-level task differences that persists throughout the context. We find cases where models do not form a global task representation at particular token sites, such as tasks involving state tracking and tasks combining distinct subtasks together. We also find nuances in the distinction between the two types of task representations, where the representational and functional roles of different tokens exhibit non-trivial overlap in certain settings.
In general, our results complicate the intuitive picture that language models smoothly and gradually refine task representations during in-context learning (ICL). Models form task representations at different levels and different rates. Across tokens and examples in the context, the same task state is not sustained, but can fade and reactivate. These high-level changes in models’ representational states unite a few observations from prior work (bakalova2025contextualize; cho2024revisiting; dong2025understanding). Key tokens prior to answer generation help aggregate in-context examples but do not all participate in prediction circuits, as the effective task representations they once generate are deactivated and reactivated later. The convergence of transferrable task representations may result from the contextualization subcircuit identified in (bakalova2025contextualize), such that later demonstrations contribute more to the test query as seen in (cho2024revisiting). Tokens seemingly outside of a core ICL circuit may nonetheless help restore task contexts to some extent, potentially contributing to the bypass mechanisms noted in (cho2024revisiting).
One trend that arose from these investigations is that language models do not seem to condense task information into local representations in all cases. We find that language models can construct token-level task representations that flexibly capture more broad task contexts and longer outputs. But in some cases, models elect to form sharp local contexts only for small task scopes and offload broader task representations across multiple tokens. This flexibility across different types of longer-generation tasks adds additional nuance to the distributed-ness of effective task vectors across multiple tokens similarly noted in dong2025understanding and tikhonov2025one. This flexibility may also relate to the success many studies have observed on extracting transferrable task representations in broad settings, such as following instructions or images (davidson2025different; huang2024multimodal; luo2024vision), or capturing information for multiple possible task outcomes (xiong2024everything).
@@
For some tasks that require more complicated computation such intermediate-state tracking, successful inference may need to rely on not only cross-token but cross-layer representations (e.g., as shown in ameisen2025circuit). In these cases, the effective restoration of the computation process may also require intervening multiple layers during the forward pass. It is also possible that by overriding token activations at intermediate layers with task vectors, the models have lost task state information formed in earlier layers. Here, we show that zero-shot recontextualization remains challenging in these tasks even through additive injection such as function vectors (todd2023function, see Appendix D). But more advanced methods that restore task states by intervening multiple layers may be more successful at restoring model task states in these cases and elicit different temporal or semantic dynamics (li2024implicit; saglam2025learning). We also show some evidence that different models may differ in the extent to which they form local or distributed representations in different tasks.
@@
An intriguing direction for future work is to study the mechanistic bases for the strong temporal and scope locality we observed in models’ in-context transferrable task representations. One possibility may be that the residual stream is more stable and easier to learn from during training. This may encourage the model to rely more on the residual stream to condense contextual task representations rather than relying on the more expensive attention operations. These learning dynamics may drive models to conform with an implicit normative consideration to not instantiate task contexts until needed and instantiate just the right scope to avoid capacity waste. Some of these features may even be exclusive to models pre-trained on natural languages (e.g. yang2025task).
Taken together, our findings offer some useful insights for interpretability and interactions with language models at large. For the science of understanding language models, comparing between more passive and more active, transferrable levels of representation offers a useful framework for probing representations of different knowledge or concepts. Practically, the variant temporal and semantic locality of different task representations means that interactions such as prompting or model steering should not assume a stable, continuous task state. Rather, effective control may require strategic re-instantiation of the fleeting representations at critical subtask boundaries.
@@
We note a few important limitations of our work. First, we primarily rely on single-token, single-layer intervention methods to extract transferrable task representations. This means that our conclusions and speculations are bounded by the effectiveness of these methods. As we discussed earlier, representations for some task contexts may be more distributed, either across tokens and/or across model layers. It would be important to confirm if similar dynamics are observed in less-constrained methods such as a multi-layer recontextualization (li2024implicit; saglam2025learning). Second, we mostly explored relatively simple tasks, including when we investigated longer-generation tasks. It’s possible that many of the dynamics we observe here would not generalize to settings with naturalistic languages, especially when the tasks are not cleanly decomposable and a single, semantically-independent task scope is hard to define.
@@ Conclusion
We investigated how the dynamics of in-context learning are reflected in the development of language models’ internal task representations. Our results suggest that language models do not smoothly refine a global task state. While general task sensitivity persists throughout the context, models construct more active, transferrable task representations in a “just-in-time” fashion. These contrasting levels of task representations provide new insights into the models’ state of inferring and performing tasks based on new evidence.
@@
We thank Yasaman Bahri, Katherine Hermann, Mike Mozer, Sjoerd van Steenkiste, Murray Shanahan, and anonymous reviewers for helpful comments on this work.
@@ Appendix A Tasks
Table 1: Simple/shorter-answer tasks. See todd2023function for more details for the first nine tasks.
@@ Appendix C Results on Qwen3 models
Appendix D Temporal and semantic locality in Function Vectors
We show that the main locality features of transferrable task representations replicate when using function vectors (todd2023function) as the extraction and recontextualization method. Function vectors are the summed activations of a set of critical attention heads that contribute to in-context task performance. Similar to task vectors, function vectors can transport task semantics across a range of tasks when additively injected onto the last token in a zero-shot prompt.
We build on the procedure outlined in todd2023function to extract function vectors from the key last “:” token in 8-shot prompts, with the following exception: we separately measure the critical attention heads contributing to a particular task and extract a per-task function vector. We use the top 16 heads for QWEN3_4B and GEMMA_V3_4B, 32 heads for QWEN3_8B, 48 heads for GEMMA_V3_12B, 64 heads for QWEN3_14B, and 80 heads for GEMMA_V3_27B, roughly corresponding to the number of critical attention heads used in similar-sized models in todd2023function. To remain comparable with the task vector extraction procedure mentioned above, we similarly sampled at most 50 queries per task to extract function vectors and search for the best intervention layer in zero-shot prompts. We then test the function vector with the best-layer intervention on the remaining queries. In the function vector case, the development set queries are sampled from cases where models correctly generate the target answer in an 8-shot prompt.
Finally, we generalize the extraction of function vectors on other tokens in the prompt to study when this kind of transferrable task representation form. We re-compute the critical attention heads and the best intervention layer per token site. Figures S15 and  S16 show replication of the main temporal and semantic scope locality of transferrable function vector representations on Qwen3 models. These results suggest that function vectors and task vectors extract similar kinds of transferrable task representations during models’ in-context learning. We unfortunately have yet to be able to induce zero-shot task recontextualization using function vectors on Gemma3 models. We suspect that Gemma3 models’ normalization scheme and the increasing magnitude of their residual streams may require scaling up the signals from function vectors when intervening on later layers, but this is a space requiring more investigation.
Figure S14: Recontextualized zero-shot task performance using function vectors.
Figure S15: Effective function vectors, like task vectors, tend to only activate on the key colon tokens.
The plot shows the zero-shot accuracies when the models are recontextualized with the per-task function vector. Accuracies are averaged across the 14 simple tasks.
Figure S16: Function vectors extracted from the key colon token similarly capture limited scope locality, especially in cases where the overall task combines semantically independent subtasks.
Visualization as in Figure 5.
Appendix E Cross-token analyses
We conduct a set of cross-token transfer and identification experiments to further understand the representational and functional roles of non-key tokens, and the overlap between identifiable and transferable task representations.
Cross-token task vector transfer We first test two variants of cross-token task vector intervention. We use representations from non-key tokens in few-shot prompts to intervene the key “:” token in zero-shot prompts (Figure S17). We also test using representations from the key “:” token in few-shot prompts to intervene different token sites in zero-shot prompts (Figure S18). For each token pair in either experiment, we re-compute the most effective layer of the cross-token intervention.
In both cases, we observe some non-trivial success of task recontextualization. Although there is notable task and model dependence, we find that representations of the “:” token after the “Q” token (prior to query presentation) can be used to help restore task contexts in zero-shot settings. We also find that in Qwen3 models, multiple non-key token sites can functionally participate in task recontextualization when given effective transferrable task representations. These results suggest a degree of shared task identity representation and transferrable task representation in some tokens. They also indicate that token sites outside of a core circuit for in-context learning (ICL) may nonetheless preserve a useful functional role when given the right information. This could potentially contribute to bypass mechanisms noted in (cho2024revisiting), where some residual ICL ability remains in the model even when ablating core ICL circuits.
Figure S17: Recontextualized zero-shot accuracy across tasks, using representations from different tokens in 8-shot prompts to intervene the last “:” token in zero-shot prompts. The x-axis indicates the token representations used to intervene zero-shot inference. Each row represents one task. We exclude samples where the correct output is simply the first word of the input query.
Figure S18: Recontextualized zero-shot accuracy across tasks, using representations from the last “:” token in 8-shot prompts to intervene different tokens in zero-shot prompts. The x-axis indicates the token site intervened during zero-shot inference. Each row represents one task. We exclude samples where the correct output is simply the first word of the input query.
Analyzing how task identity representations overlap with transferable ones at key tokens: In this section, we further analyze the relationship between the identifiable and transferable task representations. Specifically, we fit linear classifiers that extract identifiable task representations across all the non-effective format tokens—i.e., all format tokens except the 2nd ‘:’ where we find transferable task representations. We also test a version without the first ‘:’ token, given the results above. We then test how well these classifiers generalize to identifying the task representations on the held-out 2nd ‘:’ token representations from a disjoint set of prompts.
We find that there is reliable generalization (Table 4; Figure S19). In fact, we observe perfect generalization for all tasks involving a single output, as well as some of the mixed-generation tasks such as antonym + product-company. However, for many list-operations tasks (e.g. reverse_all_of_3), results are somewhat less reliable; but the classifier still achieves robust 75-80% generalization (far above chance performance of 3%, or 6% among these tasks). The performance is comparable with or without the first ‘:’ in the training set for the classifier, suggesting that this generalization is not primarily due to the partially-transferable representations on that token observed above.
Together, these results suggest that identifiable task representations persist in a relatively consistent way across both tokens that do not produce transferable task representations, and those that do produce transferable task representations. This suggest that perhaps the identifiable and transferable task representations lie in subspaces that are not fully overlapping.
To understand these results more deeply, we analyzed the relationship between the linear dimensions identified by the regression, and the dimensions of variation in the transferable task representations (Figure S20). We performed PCA over the transferable representations, and identified how strongly the regression features projected onto the top 1-20 principle components. Consistent with the results above, we found that for the same cluster of simple and non-list-operations tasks for which the regressions generalized best, the regression also projected fairly strongly onto the top principle components (with 40%-60% of the identifiable dimension projecting onto the top 20 principle components of transferable representations). By contrast, identifying the other list-operation tasks generally seemed to rely more heavily on components outside the top PCs, with only 10-25% of the identifiable dimension projecting onto the top 20 principle components of the transferable representations. Correspondingly, we find that if we project the transferable representations down to the top 20 principle components, we find strong generalization (99.1%) for the simpler tasks compared to weaker generalization for the other list-operations tasks (34.8%).
Together, these results suggest that identifiable and transferable task representations lie in partially-overlapping subspaces of the models representations, but the degree of that overlap is modulated by the task type in similar ways to our other findings. These results reinforce our suggestion that these distinct types of task representation are distributed differently across the sequence, but shed more light on their interactions and their overlap.
Table 4: Task-identifying regressions generalize from non-transferable tokens to transferable ones.
Training tokens Test tokens Task set Generalization Top-20-PCs gen.
[…]
Other tasks 1.0 1.0
[…]
Other tasks 1.0 -
Figure S19: The confusion matrix for the task-identifying regressions (when tested on generalization from non-transferable to transferable representations) shows more details of the errors—while many tasks are identified perfectly, some of the list operation tasks yield more errors. Nevertheless, generalization performance is relatively high even among the list-operations tasks.
Figure S20: Identifiable task representations partly overlap with the top principle components of transferable task representations, but the degree of this overlap is modulated by the task type.
Additional details: We performed these experiments in layer 29 of Gemma_V3_27B_PT, which was identified as the mode best layer for transferable task representations. We used 8-shot prompts, and fit the linear classifiers via $\ell^{2}$-regularized logistic regression. We also fit regressions within only the shared-vocab tasks used in some of the analyses above, and we observe qualitatively similar patterns of results in that setting.
Appendix F Inspecting subtask representations in mixed-generation tasks
We showed above that, in mixed-generation tasks, recontextualizing zero-shot inference with task vectors extracted from the key “:” was not enough to support continued generation of subsequent subtask answers. Here, we explore two hypotheses of how subtask representations may be restored through intervention. We take representations from the “:” token in 8-shot prompts and intervene at the “,” tokens prior to subtask answers in zero-shot prompts. We compare this intervention to using representations from the matching “,” tokens prior to subtask answers in 8-shot prompts to intervene the “,” tokens during zero-shot answer generation. Table 5 demonstrates the intervention success when “,” tokens prior to subtask generation are intervened with representations from matching prior “,” tokens, rather than representations from “:” tokens. The results are consistent with findings from tikhonov2025one, suggesting that the formation of transferrable subtask representations can be delayed to later format tokens, especially in cases where the subtasks are semantically independent. However, we also note that the restored subtask accuracies are still below that of single-task recontextualization. This may reflect that the full subtask representation requires a combination of task representation in “:” tokens as well as in “,” tokens.
Table 5: Interventions on comma tokens in mixed-generation tasks. The table shows the intervened zero-shot accuracy for the output immediately following the intervened “,” token versus any remaining output. Results are averaged across tasks.
Source token Model Output after intervened “,” Remaining Output
“:” Gemma_V3_4B_PT 0.004 0.000
Gemma_V3_12B_PT 0.004 0.003
Gemma_V3_27B_PT 0.003 0.014
Qwen3_4B 0.003 0.000
Qwen3_8B 0.001 0.000
Qwen3_14B 0.004 0.009
matched “,” Gemma_V3_4B_PT 0.455 0.002
Gemma_V3_12B_PT 0.295 0.001
Gemma_V3_27B_PT 0.413 0.007
Qwen3_4B 0.198 0.003
Qwen3_8B 0.293 0.011
Qwen3_14B 0.248 0.016